OpenStudy (anonymous):

Consider the functions f(x) = 3x2 and g(x) = −2x − 5. What is the value of f[g(−4)]? 9 −31 27 3

OpenStudy (anonymous):

This is asking you to compose the functions, so begin with the g function and substitute -4 in for x to get an answer. Take this answer and substitute it into x in the f function.

OpenStudy (anonymous):

Nope, you try it. I'll tell you if you're correct or not.

OpenStudy (anonymous):

ok, first try to find g(-4) you know that g(x) = −2x − 5 now just instead of x plug in -4 g(-4) = -2(-4) -5 = ? solve the equation to find ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Good. Now put 3 in for x in the f function.

OpenStudy (anonymous):

I solved it and got -31. Is that correct?

OpenStudy (anonymous):

Isn't f(x) = 3x^2???? Try again.

OpenStudy (anonymous):

f(x) = 3*x^2 f(3) = 3*(3^2) = ? solve for ?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

Nicely done.

OpenStudy (anonymous):

So is that the answer?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Can u help me with 2 more problems like this?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f(x) = 8x + 2 and g(x) = 5x - 3. What is the value of f/g (4)?